# Explanation Space: A New Perspective into Time Series Interpretability

### v1

@@ Introduction
where subscript $c$ indicates the probability output of the model for class $c$, the predicted class for sample $x$. Similar to (Höllig, Thoma, and Grimm 2023), the XAI method’s robustness is defined as follows:
$Rbt_{xai}=\frac{1}{|x|}||E(M(.),x)-E(M(.),x+\lambda\epsilon)||$ (4)
@@
Faithfulness: An explanation is faithful if the input features with highest attributions have the most effect on the prediction (Löffler et al. 2022). We measure the faithfulness using percentage of samples for which the prediction label flips when all non-negligible attributions are replaced with a default value, as follows:
@@
In literature, some studies proposed using values rather than zero as a default baseline, either for generating explanation or measuring faithfulness. For example, (Löffler et al. 2022) replaces a time step with the average value of that time step in training set (which is not a practical assumption for shift-variant time series), (Schlegel et al. 2019) picks a subsequence and reverses the time orders of the time steps, and (Parvatharaju et al. 2021) learns the most useful replacement/perturbation from the training data, etc. Such assumptions are always data type dependent. We avoid the baseline problem in all non-time domain spaces by defining them in such a way that zero has a specific meaning corresponding to certain lack of features. However, whether a certain space is suitable for a target dataset needs to be carefully considered by a practitioner familiar with that type of time series.
@@
Sparsity: In time series explanation, sparsity (also known as complexity) is often defined using Shannon entropy (Höllig, Thoma, and Grimm 2023; Vielhaben et al. 2024). Despite its wide usage, it has a few limitations. First, it does not have a predefined upper limit. In other words, for a given dataset, Shannon entropy is always between zero and a positive real number that depends on the length of the time series. Because the upper limit is not fixed and predefined, it is hard for an XAI user to internalize how complex an explanation is in an absolute sense. Another problem with Shannon entropy is its length-dependent accumulation of extremely small non-zero attributions. A motivating example of why the length matters is shown in Figure 4. Since the length of the input in time domain is twice the length in frequency domain and because the attribution in time domain contains many small non-zero values, all non-zero values accumulate resulting in the higher Shannon entropy in time domain. This is in stark contrast to what is visually presented. Our formulation (in equation (6)), however, indicates better sparsity in time domain since it takes the length into account.
Last, ignoring the length may have other consequences. For example, imagine two different input domains, $X_{1}$ and $X_{2}$, where the length of the first one is larger than the second one, i.e. $l_{1}>>l_{2}$. If an XAI method assigns high attribution to two input features and zero to others in both input domains, Shannon entropy of the two explanations is equal. However, picking out two input features in an space where the length of the input is larger is more significant. Although this was not a problem in previous studies due to the fix length of the time domain, in our evaluation we need a metric sensitive to the input length because some explanation spaces, such as frequency and time/frequency, have different lengths than the corresponding time domain. We note that the choice of an explanation space and what metric is appropriate is not trivial and depends on the nature of the application. Hence, a practitioner should choose a suitable option accordingly.
Here, we introduce four desirable properties for a sparsity metric and then propose a new simple metric that satisfies all properties. A sparsity metric, $Spr(.)$, should satisfy:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
[…]
where $\beta>1$ is a constant hyper-parameter to spread out the metric more uniformly. It is easy to show that it satisfies all above properties. See Appendix for more details.
@@
Table 2: ResNet: The first number is the faithfulness and the number in the parenthesis is the sparsity metric.
Sparsity on Different Spaces
In Table 2, we report sparsity along the faithfulness success percentage to roll out sparse but non-faithful explanations. Here, the first number is the success percentage and the number in the parenthesis is the sparsity. We exclude the sparsity for entries with lower than $50\%$ success percentage. For most datasets, LIME only picks a few input features (regardless of the input space) and produces a highly sparse explanation. However, most explanations turn out to be irrelevant in terms of faithfulness. Note that we use default hyper-parameters when possible for all XAI methods. It is possible to make some methods better by tuning the parameters. For example, by grouping each $m$ consecutive time steps as a single feature, LIME can generate more faithful explanation. However, such a hyper-parameter search is not only beyond the scope of this paper, but also unnecessary for the purpose of this paper. We emphasize that our goal is to show that the same XAI method can be used on different explanation spaces and generate different quality explanation.
Dataset Domain Classifier XAI Robustness (×$\times 10^{-4}$)
AudioMNIST Time 73.7 3.7 13.5 4.9 5.5 3.4 14.4 6.3 4.0 24.3
Freq. 85.5 7.0 9.4 2.0 8.3 3.0 10.6 6.8 10.0 11.6
Time/Freq. 18.9 0.6 5.7 0.3 1.1 0.5 11.3 3.1 0.7 5.5
Difference 46.8 4.4 13.9 16.0 4.6 5.3 14.4 6.9 3.3 37.4
Min Zero 78.3 18.3 23.0 5.0 24.5 16.1 23.8 6.7 0.8 24.6
FordA Time 2.5 12.6 81.6 13.5 19.3 13.2 108.8 51.7 3.6 27.0
Freq. 2.6 1.1 34.8 0.8 1.9 1.0 79.1 39.5 2.2 2.6
Time/Freq. 4.2 1.7 24.7 1.8 3.0 1.1 57.6 23.5 1.0 3.8
Difference 19.7 20.2 92.3 38.3 44.2 22.6 66.7 56.2 4.0 96.6
Min Zero 2.6 17.9 81.9 13.7 27.0 20.8 115.0 53.2 0.2 27.2
GunPoint Time 15.2 129.5 174.9 43.7 157.8 160.2 201.9 177.2 17.9 152.7
Freq. 7.0 6.8 61.3 2.0 23.1 13.6 196.5 163.2 5.9 26.8
Difference 3.6 5.6 175.5 12.2 19.3 6.8 217.6 178.1 9.5 43.3
Min Zero 16.0 55.7 171.2 43.1 93.6 77.3 207.1 175.2 12.3 152.2
ECG5000 Time 4.2 35.5 160.5 44.8 72.1 35.7 237.5 188.4 22.6 94.0
Freq. 4.7 8.2 99.5 5.3 24.2 10.1 201.0 186.5 20.2 34.5
Difference 6.0 2.7 63.4 31.0 9.2 10.0 214.3 185.6 4.8 60.3
Min Zero 5.0 69.8 188.6 51.4 100.8 76.9 241.4 191.1 4.9 102.8
LargeKit.App. Time 5.1 1.7 30.1 2.5 2.9 1.5 87.3 36.2 3.3 13.8
Freq. 22.4 14.7 31.3 7.5 19.7 5.7 56.8 37.3 3.2 24.5
Time/Freq. 3.0 1.2 16.0 0.9 2.2 1.2 45.6 13.1 0.6 2.3
Difference 157.7 6.4 26.1 66.9 5.0 6.1 92.1 39.9 1.2 81.6
Min Zero 5.0 1.4 30.3 2.5 2.8 1.5 54.4 35.8 2.6 13.8
Table 3: Robustness (ResNet)
As shown in Table 2, different XAI methods may produce the most sparse explanation in different spaces. So, the sparsity is not always consistent. Nevertheless, the sparsity pattern aligns with our intuition when we proposed explanation spaces. For example, AudioMNIST, containing audio signals, is the most sparse in time/frequency domain in 6 out of 8 faithful methods and the two others are frequency spaces. FordA, a dataset of engine noises, is most sparse in frequency domain in 6 out of 8 faithful methods, as it was shown in Figure 1a and 1b. As we speculated, both ECGFiveDays and ECG5000 are sparsest in difference space for most XAI methods.  [^6] Note that we exclude experiment with time/frequency domain in datasets where the number of time steps is small because it was impossible to get a better resolution than the frequency domain. The last three datasets, LargeKitchenAppliance, ElectricDevices, and Earthquakes, are examples of datasets where the original zero value (before normalization) correspond to the lack of feature and, hence, min zero space is the most sparse space across most XAI methods. Overall, the only exception we encounter was the GunPoint dataset where we initially hypothesized that the difference space is the most appropriate due to the nature of the dataset, containing hand movement. However, by comparing time and frequency domains (as depicted in Figure 2e and 2f), it is evident that only the first few frequency components are non-negligible and others are relatively small. Hence, we speculate that for datasets containing only low frequencies, any XAI method is likely to be more sparse in the frequency space. Due to the lack of space, the results for InceptionTime is provided in Appendix. Overall, we observe the same pattern.
Robustness on Different Spaces
The results of robustness for ResNet model is shown in Table 3. Due to the lack of space, we only show one dataset for each dataset type. See Appendix for results on other datasets and also InceptionTime model. Unlike sparsity metric, there is no strong correlation between dataset types and robustness. However, there are a few observations worth considering: First, classifier robustness does not necessarily match the XAI robustness. For example, the time domain is the most robust space in ECG5000 while the XAI robustness on time domain are among the worst. Interestingly, whenever time/frequency is provided for a dataset, XAI methods on this space are among the most robust ones (20 out of 27). Unlike the classifier robustness where the min zero space and time domain perform similarly, their XAI robustness can be vastly different due to the way they handle the baseline. Overall, no clear indication can be given on the best explanation space purely based on the robustness.
Discussion
Unlike computer vision, time series data is an umbrella term covering vastly different applications with widely different characteristics. This includes stock market data, EEG signal, audio signals, accelerometer data, as well as non-temporal data, such as spectrogram (e.g. Beef and Coffee dataset in UCR repository) or skeleton(contour)-based descriptors (e.g. ArrowHead or Fish datasets). Spplication-specific experts often rely on domain-specific knowledge for time series analysis that is absent and disconnected from the computer science community. This leads to the proliferation of classification and explanation methods that do not take application specific nuances into account (e.g. the usage of different spaces/domains.). While the application-blind approach has often worked largely for classifications due to the power DL, it faces significant challenges when it comes to explanation.
In this paper, we bridge the gap between application-blind approaches in computer science community and the application-specific approaches used by domain experts. Our method allows computer scientists to continue proposing application-blind XAI methods without worrying about the explanation space. Interdisciplinary researchers can now focus on proposing new mappings to project known spaces to desired, yet ignored, spaces, to facilitate the usage of existing XAI. Finally, our framework allows domain experts to easily explore other spaces and investigate if there exists a better space for explanation, similar to what we have observed from GunPoint dataset where frequency domain tends to provide more sparser explanation rather than the hypothesized difference space. This can also help practitioners to investigate what features a model uses. As shown in (Vielhaben et al. 2024), a model that exploits average value of the entire time series is hard to catch in time domain. However, it can be easily identified by checking the first frequency component of the frequency space.
Conclusion
In this paper, we introduce a wrapping process that allows a model trained on one domain to operate on other domains and, thus, allows existing XAI methods to explain the model in different input domains. We suggest four new spaces. Moreover, we introduce a new sparsity metric satisfying a few desirable properties. Using the proposed metric, we show that different explanation spaces can be useful for different types of time series consistent with domain knowledge and intuition. This simple yet effective technique allows a practitioner familiar with a time series in hand to easily pick the appropriate explanation space and explain the trained classifier using any existing XAI methods.
Appendix A Appendices
Explanation Space Implementation
The representation space function, $F(x)$, is used outside a neural network and it can be construed as a pre-processing stage (with respect to the explanation generation stage at inference time, not with respect to training stage). Hence, the target domain data, $z=F(x)$, can be obtained by implementing the transformation in any arbitrary way. However, $F^{-1}(z)$ function is embedded into the $M^{\prime}$ neural network. Hence, for compatibility with gradient-based methods, we either need to use the existing functions in deep learning packages or implement the function using existing layers. Fortunately, most packages, like PyTorch, has the implementation of FFT, STFT (through torchaudio.transforms.Spectrogram), and their inverse. We implement inverse mapping function for min zero space and difference space using a single fully connected layer with a pre-defined weight matrix, as follows:
Min Zero Space Implementation: The inverse mapping of the min zero space can be written as follows:
$F^{-1}_{min\_zero}(\{x_{1}-x_{min},...,x_{N}-x_{min},x_{min}\})\rightarrow\\ \{x_{1},x_{2},...,x_{N}\}$ (7)
In terms of a fully connected layer, it can be implemented as a layer that takes a vector of size $N+1$ as input and output a vector of size $N$, where the $(N+1)^{th}$ element is added to all other elements in the input vector and then it is dropped from the output. It can be easily shown that an Identity matrix of size $N\times N$ concatenated to an all-one matrix of size $1\times N$ can obtain the reverse mapping as a linear layer as follows:
$\begin{bmatrix}x_{1}-x_{min}\\ x_{2}-x_{min}\\ ...\\ x_{N-1}-x_{min}\\ x_{N}-x_{min}\\ x_{min}\\ \end{bmatrix}_{N+1\times 1}^{T}\begin{bmatrix}1&0&...&0&0\\ 0&1&...&0&0\\ ...&...&...&...&...\\ 0&0&...&1&0\\ 0&0&...&0&1\\ 1&1&...&1&1\\ \end{bmatrix}_{N+1\times N}\\ =\begin{bmatrix}x_{1}\\ x_{2}\\ ...\\ x_{N-1}\\ x_{N}\\ \end{bmatrix}_{N\times 1}^{T}$ (8)
Difference Space Implementation: The inverse mapping of the difference space can be written as follows:
$F^{-1}_{diff}(\{x_{1},x_{2}-x_{1},x_{3}-x_{2},...,x_{N}-x_{N-1}\})\rightarrow\\ \{x_{1},x_{2},...,x_{N}\}$ (9)
It can be easily shown that the $i^{th}$ elements of the output can be obtained by adding all previous input elements up to $i^{th}$ element. As a result, all $1$ to $i-1$ elements are cancel out and only the $x_{i}$ remains which constructs the inverse operation. In linear layer with matrix multiplication operation, it can be easily implemented by a weight matrix of upper triangular form where every non-zero element is exactly one, as follow:
$\begin{bmatrix}x_{1}\\ x_{2}-x_{1}\\ ...\\ x_{N-1}-x_{N-2}\\ x_{N}-x_{N-1}\\ \end{bmatrix}_{N\times 1}^{T}\begin{bmatrix}1&1&...&1&1\\ 0&1&...&1&1\\ ...&...&...&...&...\\ 0&0&...&1&1\\ 0&0&...&0&1\\ \end{bmatrix}_{N\times N}\\ =\begin{bmatrix}x_{1}\\ x_{2}\\ ...\\ x_{N-1}\\ x_{N}\\ \end{bmatrix}_{N\times 1}^{T}$ (10)
What Is $\beta$ In Sparsity?
In the sparsity metric, $\beta>1$ is a constant hyper-parameter to spread out the metric more uniformly. Empirically, for majority of datasets/XAIs, the attribution is relatively sparse and it is rarely uniformly distributed. Consequently, in most cases, all methods give a value between 0.90 and 1 which makes it hard to compare one another. Hence, choosing $\beta>1$ allows us to better spread out the values and compare XAI Methods.
Sparsity/Faithfulness Results on InceptionTime
The results of the InceptionTime model is presented in Table 4. The results are similar to ResNet model. For example, AudioMNIST is mostly sparse on time/frequency and frequency domains. FordA is mostly sparse on frequency domain and so on.
@@ Appendix A Appendices
Table 4: InceptionTime: The first number is the faithfulness success percentage (SP), $Faith$, and the number in the parenthesis is the sparsity metric.
@@
ECGFiveDays Time 7.1 13.0 122.5 22.9 18.6 11.2 242.9 188.2 13.5 98.6
Freq. 15.6 35.4 111.1 10.5 44.6 16.5 182.5 190.6 11.4 49.7
Difference 19.0 36.9 92.7 37.3 43.4 24.0 218.5 199.9 13.5 188.7
Min Zero 10.4 94.3 152.7 29.0 111.4 86.5 253.0 195.0 12.2 114.1
ElectricDev. Time 28.9 39.7 189.7 14.1 156.0 43.5 309.6 223.7 31.0 237.2
Freq. 217.5 127.2 190.5 32.5 160.1 69.0 253.7 220.0 96.9 173.9
Difference 111.6 70.8 140.7 257.5 96.4 59.3 276.4 275.3 42.8 306.6
Min Zero 29.4 12.0 163.4 14.1 38.4 13.5 284.8 236.2 12.7 236.1
Earthquakes Time 1.3 1.4 56.2 1.7 4.1 2.2 59.7 39.5 2.0 6.9
Freq. 2.6 3.7 67.6 2.9 8.0 4.2 76.5 34.8 5.7 10.7
Time/Freq. 0.3 1.9 34.1 1.3 4.3 1.7 53.5 35.9 1.0 4.5
Difference 35.0 27.3 83.2 35.1 41.1 38.2 80.4 55.2 8.0 97.2
Min Zero 1.3 1.3 57.9 1.7 3.8 2.1 65.1 43.1 1.8 6.9
Table 5: Robustness (ResNet)
Robustness Results
The rest of the robustness results on ResNet is shown in Table 5. The results of classifier robustness and XAI robustness is presented in Table 6. The overall patterns are similar to ResNet.
@@
FordA Time 1.7 nan 81.7 14.7 15.5 13.2 93.6 52.2 3.3 20.6
Freq. 2.2 nan 37.3 0.9 2.5 1.0 78.2 46.2 2.2 3.5
Time/Freq. 7.5 nan 23.3 3.9 2.9 1.2 55.8 22.8 0.9 3.2
Difference 13.2 nan 82.9 82.9 45.9 26.1 78.9 57.0 4.1 97.2
Min Zero 1.9 nan 81.4 14.9 18.3 16.5 118.0 53.7 0.2 20.8
@@
ElectricDev. Time 292.5 45.1 213.4 137.3 127.8 101.5 291.3 217.2 73.0 230.1
Freq. 375.0 56.3 187.1 91.8 153.7 72.0 245.9 203.2 105.8 174.7
Difference 291.2 56.6 123.1 51.4 114.0 58.0 282.3 265.5 72.0 310.8
Min Zero 289.8 39.0 173.3 136.0 35.6 22.9 284.2 246.1 23.2 228.4
Earthquakes Time 1.0 6.8 63.5 13.4 6.4 4.5 48.9 51.6 1.8 22.9
Freq. 1.7 14.8 67.9 9.6 9.7 3.8 54.7 51.4 7.8 13.3
Time/Freq. 1.3 7.7 38.0 4.0 4.8 2.0 47.9 20.1 1.3 5.3
Difference 39.6 21.9 70.0 33.2 48.5 31.5 82.0 56.5 8.9 115.0
Min Zero 0.9 6.9 65.4 13.3 4.1 2.5 47.4 51.7 1.5 22.9
Table 6: Robustness (InceptionTime)
Reproducibility Checklist
This paper:
  • Includes a conceptual outline and/or pseudocode description of AI methods introduced (yes/partial/no/NA)? Yes.
  • Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results (yes/no) Yes.
  • Provides well marked pedagogical references for less-familiare readers to gain background necessary to replicate the paper (yes/no) Yes.
Does this paper make theoretical contributions? (yes/no) No.
Does this paper rely on one or more datasets? (yes/no) Yes.
If yes, please complete the list below.
  • A motivation is given for why the experiments are conducted on the selected datasets (yes/partial/no/NA) Yes.
  • All novel datasets introduced in this paper are included in a data appendix. (yes/partial/no/NA) No. They are publicly available
  • All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes/partial/no/NA) NA.
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations. (yes/no/NA) Yes.
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. (yes/partial/no/NA) Yes.
  • All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisfying. (yes/partial/no/NA) NA.
Does this paper include computational experiments? (yes/no) Yes.
If yes, please complete the list below.
  • Any code required for pre-processing data is included in the appendix. (yes/partial/no). Yes.
  • All source code required for conducting and analyzing the experiments is included in a code appendix. (yes/partial/no) Yes.
  • All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes/partial/no) Yes.
  • All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from (yes/partial/no) Partial.
  • If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. (yes/partial/no/NA) Yes. All experiments are started by setting the randomness seed to zero.
  • This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks. (yes/partial/no) No.
  • This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. (yes/partial/no) Yes.
  • This paper states the number of algorithm runs used to compute each reported result. (yes/no) Yes.
  • Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. (yes/no) No.
  • The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). (yes/partial/no) No.
  • This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. (yes/partial/no/NA) Partial.
  • This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. (yes/partial/no/NA) Partial.

### v3

@@ IV-A Evaluation Criteria
where subscript $c$ indicates the probability output of the model for class $c$, the predicted class for sample $x$. Similar to [18], the XAI method’s robustness is defined as follows:
$Rbt_{xai}=\frac{1}{|x|}||E(M(\cdot),x)-E(M(\cdot),x+\lambda\epsilon)||$ (4)
@@
Faithfulness: An explanation is faithful if the input features with highest attributions have the most effect on the prediction [30]. We measure the faithfulness using percentage of samples for which the prediction label flips when all non-negligible attributions are replaced with a default value, as follows:
@@
In the literature, some studies proposed using values rather than zero as a default baseline, either for generating explanation or measuring faithfulness. For example, [30] replaces a time step with the average value of that time step in training set (which is not a practical assumption for shift-variant time series), [31] picks a subsequence and reverses the time orders of the time steps, and [32] learns the most useful replacement/perturbation from the training data, etc. Such assumptions are always data type dependent. We avoid the baseline problem in all non-time domain spaces by defining them in such a way that zero has a specific meaning corresponding to certain lack of features. However, whether a certain space is suitable for a target dataset needs to be carefully considered by a practitioner familiar with that type of time series.
@@
TABLE II: ResNet: The first number is the faithfulness and the number in the parenthesis is the sparsity metric.
@@
TABLE III: InceptionTime: he first number is the faithfulness and the number in the parenthesis is the sparsity metric.
Sparsity: In the literature, sparsity (also known as complexity) is measured using Shannon entropy [18] [15]. Shannon entropy has a few limitations: First, it does not have a predefined upper limit. In other words, for a given dataset, Shannon entropy is always between zero and a positive real number that depends on the length of the time series. Because the upper limit is not fixed and predefined, it is hard for an XAI user to internalize how complex an explanation is in an absolute sense. Another problem with Shannon entropy is its length-dependent accumulation of extremely small non-zero attributions. A motivating example of why the length matters is shown in Figure 5. Since the length of the input in time domain is twice the length in frequency domain and because the attribution in time domain contains many small non-zero values, all non-zero values accumulate resulting in the higher Shannon entropy in time domain. This is in stark contrast to what is visually presented. Our formulation (in equation (6)), however, indicates better sparsity since it takes the length into account.
Last, ignoring the length may have other consequences. For example, imagine two different input domains, $X_{1}$ and $X_{2}$, where the length of the first one is larger than the second one, i.e. $l_{1}>>l_{2}$. If an XAI method assigns high attribution to two input features and zero to others in both input domains, Shannon entropy of the two explanations is equal. However, picking out two input features in an space with twice the length is more significant. Unlike previous studies, we need our metric to be sensitive to the input space length so that we can compare explanations in different spaces. We note that the choice of an explanation space and which metric is appropriate is not trivial and depends on the nature of the application. Hence, a practitioner should choose a suitable option accordingly.
Here, we introduce four desirable properties for a sparsity metric and then propose a new simple metric that satisfies all properties. A sparsity metric, $Spr(\cdot)$, should satisfy:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
[…]
where $\beta>1$ is a constant hyper-parameter to spread out the metric more uniformly. Empirically, for majority of datasets/XAIs, the attribution is relatively sparse and it is rarely uniformly distributed. Consequently, in most cases, all methods give a value between 0.90 and 1 which makes it hard to compare one another. Hence, choosing $\beta>1$ allows us to better spread out the values and compare XAI Methods. It is easy to show that it satisfies all above properties.
@@
AudioMNIST Time 73.7 3.7 13.5 4.9 5.5 3.4 14.4 6.3 4.0 24.3
Freq. 85.5 7.0 9.4 2.0 8.3 3.0 10.6 6.8 10.0 11.6
Time/Freq. 18.9 0.6 5.7 0.3 1.1 0.5 11.3 3.1 0.7 5.5
Difference 46.8 4.4 13.9 16.0 4.6 5.3 14.4 6.9 3.3 37.4
Min Zero 78.3 18.3 23.0 5.0 24.5 16.1 23.8 6.7 0.8 24.6
FordA Time 2.5 12.6 81.6 13.5 19.3 13.2 108.8 51.7 3.6 27.0
Freq. 2.6 1.1 34.8 0.8 1.9 1.0 79.1 39.5 2.2 2.6
Time/Freq. 4.2 1.7 24.7 1.8 3.0 1.1 57.6 23.5 1.0 3.8
Difference 19.7 20.2 92.3 38.3 44.2 22.6 66.7 56.2 4.0 96.6
Min Zero 2.6 17.9 81.9 13.7 27.0 20.8 115.0 53.2 0.2 27.2
GunPoint Time 15.2 129.5 174.9 43.7 157.8 160.2 201.9 177.2 17.9 152.7
Freq. 7.0 6.8 61.3 2.0 23.1 13.6 196.5 163.2 5.9 26.8
Difference 3.6 5.6 175.5 12.2 19.3 6.8 217.6 178.1 9.5 43.3
Min Zero 16.0 55.7 171.2 43.1 93.6 77.3 207.1 175.2 12.3 152.2
ECG5000 Time 4.2 35.5 160.5 44.8 72.1 35.7 237.5 188.4 22.6 94.0
Freq. 4.7 8.2 99.5 5.3 24.2 10.1 201.0 186.5 20.2 34.5
Difference 6.0 2.7 63.4 31.0 9.2 10.0 214.3 185.6 4.8 60.3
Min Zero 5.0 69.8 188.6 51.4 100.8 76.9 241.4 191.1 4.9 102.8
LargeKit.App. Time 5.1 1.7 30.1 2.5 2.9 1.5 87.3 36.2 3.3 13.8
Freq. 22.4 14.7 31.3 7.5 19.7 5.7 56.8 37.3 3.2 24.5
Time/Freq. 3.0 1.2 16.0 0.9 2.2 1.2 45.6 13.1 0.6 2.3
Difference 157.7 6.4 26.1 66.9 5.0 6.1 92.1 39.9 1.2 81.6
Min Zero 5.0 1.4 30.3 2.5 2.8 1.5 54.4 35.8 2.6 13.8
TABLE IV: Robustness (ResNet)
@@
FordA Time 1.7 10.2 81.7 14.7 15.5 13.2 93.6 52.2 3.3 20.6
Freq. 2.2 0.9 37.3 0.9 2.5 1.0 78.2 46.2 2.2 3.5
Time/Freq. 7.5 1.5 23.3 3.9 2.9 1.2 55.8 22.8 0.9 3.2
Difference 13.2 15.2 82.9 82.9 45.9 26.1 78.9 57.0 4.1 97.2
Min Zero 1.9 12.1 81.4 14.9 18.3 16.5 118.0 53.7 0.2 20.8
@@
TABLE V: Robustness (InceptionTime)
IV-B Sparsity on Different Spaces
In Table II and III, we report sparsity along the faithfulness success percentage to roll out sparse but non-faithful explanations. Here, the first number is the success percentage and the number in the parenthesis is the sparsity. We exclude the sparsity for entries with lower than $50\%$ success percentage. For most datasets, LIME only picks a few input features (regardless of the input space) and produces a highly sparse explanation. However, most explanations turn out to be irrelevant in terms of faithfulness. Note that we use default hyper-parameters when possible for all XAI methods. It is possible to make some methods better by tuning the parameters. However, such a hyper-parameter search is not only beyond the scope of this paper, but also unnecessary for the purpose of this paper. We emphasize that our goal is to show that the same XAI method can be used on different explanation spaces and generate different quality explanation.
As shown in Table II, XAI methods may disagree on the space in which they produce the sparsest explanation. Nevertheless, the sparsity pattern aligns with our intuition when we proposed explanation spaces. For example, AudioMNIST, containing audio signals, is the most sparse in the time/frequency domain in 6 out of 8 faithful methods and the two others are frequency spaces. FordA, a dataset of engine noises, is most sparse in frequency domain in 6 out of 8 faithful methods, as it was shown in Figure 1a and 1b. As we speculated, both ECGFiveDays and ECG5000 are sparsest in difference space for most XAI methods [^6] Note that we exclude experiment with time/frequency domain in datasets where the number of time steps is small because it was impossible to get a better resolution than the frequency domain.. The last three datasets, LargeKitchenAppliance, ElectricDevices, and Earthquakes, are examples of datasets where the original zero value (before normalization) corresponds to the lack of feature and, hence, min zero space is the most sparse space across most XAI methods. Overall, the only exception was the GunPoint where we initially hypothesized that the difference space is the most appropriate due to the nature of the dataset, containing hand movement. However, by comparing time and frequency domains (as depicted in Figure 2e and 2f), it is evident that only the first few frequency components are non-negligible and others are negligible. Hence, we speculate that for datasets containing only low frequencies, sparsity metric is smaller for frequency domain. Overall, we have observed similar pattern in InceptionTime results shown in Table III.
IV-C Robustness on Different Spaces
The robustness results of ResNet model is shown in Table IV. Due to the lack of space, we show only one dataset for each dataset type. Unlike sparsity metric, there is no strong correlation between dataset types and robustness. However, there are a few observations worth considering: First, classifier robustness does not necessarily match the XAI robustness. For example, the time domain is the most robust space in ECG5000 while the XAI robustness on time domain is among the worst. Interestingly, whenever time/frequency is provided for a dataset, XAI methods on this space are among the most robust ones (20 out of 27). Unlike the classifier robustness where the min zero space and time domain perform similarly, their XAI robustness can be vastly different due to the way they handle the baseline. Overall, no clear indication can be given on the best explanation space purely based on the robustness. Results of InceptionTime reveal similar pattern, shown in Table V.
(a) Original Time Domain
(b) Decomposed component 1
(c) Decomposed component 2
(d) Decomposed component 3
Figure 6: DeepLIFT attribution method on MIMIC Performance dataset.
V Decomposition Space
To decompose a time series instance, we use singular-spectrum analysis (SSA) implementation provided in Kaggle777 https://www.kaggle.com/code/jdarcy/introducing-ssa-for-time-series-decomposition. Unfortunately, none of the datasets in the UCR repository is suitable for decomposition since they are often short with no trends or seasonality. To see the usefulness of the decomposition space, we use the MIMIC performance dataset [33] and US hourly climate dataset [34].
(a) Original Time Domain
(b) Decomposed component 1
(c) Decomposed component 2
(d) Decomposed component 3
Figure 7: Input $\times$ Gradient attribution method on US hourly dataset.
The MIMIC performance dataset contains PPG, ECG, and respiration time series and the task is to identify adults versus neonates. Here, we only use the ECG time series. We subsample each subjects’ ECG records such that each sample has a fixed length of 1000 time steps. In each sample, there are on average 25 heartbeats. Hence, the samples are long enough to capture trends or cycles. US hourly climate dataset is a public dataset provided by the National Centers of Environmental Information. It contains hourly temperature and wind speed recorded at thousands of stations across united state. Here, we use the Los Angeles airport station (USW00023174) data. We use the first $90\%$ of the time series as a training set and the last $10\%$ for validation. The length of each sample is 500 which almost covers three weeks of data. The task is to predict whether the next time step of the temperature increases or decreases. We train the ResNet50 model on the original times series (non-decomposed ones). The validation accuracy on the MIMIC and US hourly datasets was $97.93\%$ and $96.37\%$, respectively.
Figure 6 shows the attribution of the original time series versus the decomposition space components for the MIMIC performance dataset. Decomposition component 1 mostly captures the trend plus a small component of the QRS heartbeat complex. From the decomposed component 2 that has the most cyclic component of the QRS complex, it is easy to see that the model relies mainly on this cyclic component rather than the overall trend or the remaining part (the last component). This interpretation is not clearly accessible by looking at the original time series because the components are all mixed together.
Figure 7 demonstrates the attribution for the US hourly temperature dataset. Looking at the attribution in the original time domain, it is clear that the model relies mainly on the last few time points, which is reasonable for forecasting the next time step. However, it is difficult to understand what component has the most effect. In time series analysis, decomposition is often used to separate trends and cycles from the remaining part. Our approach allows different components to be studied separately. In this example, the decomposition space clearly shows that the main cycle, component 1, with largest amplitude is mostly used by the model, while the remaining two components have little effect on the model performance.
VI Discussion
In this paper, we bridge the gap between application-blind approaches in computer science community and the application-specific approaches used by domain experts. Our method allows computer scientists to continue proposing application-blind XAI methods without worrying about the explanation space. Interdisciplinary researchers can now focus on proposing new mappings to project known spaces to desired, yet ignored, spaces, to facilitate the usage of existing XAI. Finally, our framework allows domain experts to easily explore other spaces and investigate if there exists a better space for explanation. This can also help practitioners to investigate what features a model uses. As shown in [15], a model that exploits average value of the entire time series is hard to catch in time domain. However, it can be easily identified by checking the first frequency component of the frequency space.